# Causal machine learning for heterogeneous treatment effects in the presence of missing outcome data

### v1

@@ Appendix D Derivation of the EIF used within the mDR-learner and mEP-learner
$\|\hat{\theta}-\theta^{*}\|^{2}\leq$ $\frac{1}{\frac{\lambda}{2}-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]+$
$\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|^{4}+\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\mu}^{0}-\mu^{0}_{0}\|^{2}+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|^{4})$
@@
$\|\hat{\theta}-\theta^{*}\|^{2}\leq$ $\frac{1}{\frac{\lambda}{2}-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})+\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|\|\hat{\mu}^{0}-\mu^{0}_{0}\|+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|^{4})$
@@
$L(\hat{\theta},g_{0})-L(\theta^{*},g_{0})\leq$ $D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]+\frac{\beta_{1}}{2}\|\hat{\theta}-\theta^{*}\|^{2}$
$\leq$ $\frac{\beta_{1}}{\frac{\lambda}{2}-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})+\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|\|\hat{\mu}^{0}-\mu^{0}_{0}\|+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|^{4})-$
@@
$L(\hat{\theta},g_{0})-L(\theta^{*},g_{0})\leq$ $\frac{\beta_{1}}{\frac{\lambda}{2}-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})+\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|\|\hat{\mu}^{0}-\mu^{0}_{0}\|+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|^{4})$
@@
Appendix G Simulation study DGP specifications
@@
$Y(0)|Z\sim\mu_{0}(Z)+\epsilon,~{}~{}~{}Y(1)|Z\sim Y(0)+\theta(Z),~{}~{}~{}\epsilon|Z\sim\mathcal{N}(0,1)$
@@
In each DGP, 6 covariates were created, each drawn from a uniform distribution ranging from -1 to 1. The propensity score, $\pi(Z)$, used within each DGP is defined as $\pi(Z)=expit(\sum_{\{1,2,3\}}(Z_{i}^{2}-sin(3*Z_{i}))/1.5)$, and each DGP uses a combination of the following censoring, unexposed outcome and CATE functions. The complex unexposed outcome function was defined as $\mu^{0}(Z)=\sum_{\{3,4\}}(Z_{i}/2+4\mathbbm{1}\{Z_{i}>0.5\}+sin(4Z_{i}))/2$, while the simple unexposed outcome function was defined as $\mu^{0}(Z)=Z_{3}+Z_{4}$. The complex CATE was defined as $\theta(X)=\mathbbm{1}\{|Z_{3}|>0.5\}+\mathbbm{1}\{|Z_{4}|>0.5\}$, while the simple CATE was defined as $\theta(X)=0$. Then two censoring functions $G(A,Z)=expit(3-2.75A\sum_{\{4,5\}}\mathbbm{1}(|Z_{i}|>0.5))$ and $G(Z)=expit(3-2.75\sum_{\{4,5\}}\mathbbm{1}(|Z_{i}|>0.5))$ were defined, with the first imposing strong censoring only on the exposed arm, and the second imposing strong censoring evenly across both arms.
@@ 2 Complex Complex (Does not depend on A) Simple Complex
3 Complex Complex (Does not depend on A) Complex Simple
Appendix H Simulation study - Super Learner algorithm inclusion
Table 3: Super learner algorithm libraries
@@ Appendix H Simulation study - Super Learner algorithm inclusion
Appendix I Root mean square median error (RMSME)
As the DR-learner and mDR-learner use IPTWs/IPCWs, extreme propensity score/censoring probability estimates can lead to large, highly bias CATE estimates for certain individuals. Within the simulation study, this only occurs within a small subset of the simulations and consequently, the distribution of RMSE estimates ends up skewed. For this reason, a metric such as the mean RMSE can poorly represent the performance of these estimators.
Alternatively, we compare the estimators within the simulation study using the root mean square median error (RMSME). We calculate this metric by identifying the error in individual $j$’s CATE estimate for simulation $i$, $\epsilon_{ij}$. We then take the median of these errors across the simulations, $\epsilon_{j}^{med}=Median(\epsilon_{i,j})$. By doing so, extreme CATE estimates for each individual will be removed and we can then estimate the root mean squared error (RMSE) using the median errors. This metric has a similar principle to the median CATE implementation option that was outlined in Section 3.3, reducing the estimators sensitivity to extreme CATE estimates and providing a metric that ignores outlier estimates.
@@ Appendix I Root mean square median error (RMSME)
Appendix J Additional simulation study results
@@
Data Sample size mDR mEP DR EP T
Generating Available Imputation Available Imputation Available Imputation
Process Case Case Case
1 400 0.414 0.462 0.437 0.483 0.467 0.534 0.385 0.478
800 0.315 0.370 0.347 0.383 0.374 0.419 0.318 0.376
1600 0.248 0.302 0.279 0.307 0.305 0.343 0.258 0.309
3200 0.189 0.234 0.210 0.235 0.236 0.272 0.207 0.249
2 400 0.424 0.460 0.423 0.524 0.461 0.593 0.379 0.546
800 0.337 0.372 0.338 0.419 0.372 0.474 0.317 0.432
1600 0.273 0.306 0.269 0.324 0.303 0.374 0.255 0.340
3200 0.214 0.237 0.202 0.252 0.237 0.301 0.206 0.278
3 400 0.366 0.386 0.439 0.331 0.429 0.257 0.486 0.299
800 0.274 0.308 0.349 0.273 0.346 0.220 0.417 0.260
1600 0.209 0.264 0.307 0.252 0.300 0.205 0.346 0.253
3200 0.177 0.228 0.284 0.223 0.256 0.174 0.300 0.216
Figure 6: Mean RMSE for mDR-learner, mEP-learner, DR-learner, EP-learner and T-learner in three DGPs plotted by training sample size. Figures in the left column compare the mDR-learner and mEP-learner to the available case versions of the DR-learner, EP-learner and T-learner in DGP 1, 2 and 3 respectively. Figures in the right column compare the mDR-learner and mEP-learner to the DR-learner, EP-learner and T-learner when used in combination with an outcome imputation model in DGP 1, 2 and 3 respectively.
@@
Data Sample size mDR mEP DR EP T
Generating Available Imputation Available Imputation Available Imputation
Process Case Case Case
1 400 0.549 0.544 0.584 0.575 0.533 0.593 0.571 0.570
800 0.457 0.456 0.578 0.479 0.442 0.496 0.464 0.475
1600 0.547 0.371 0.528 0.419 0.369 0.381 0.415 0.391
3200 0.513 0.282 0.405 0.399 0.285 0.293 0.466 0.292
2 400 0.632 0.545 0.529 0.650 0.534 0.659 0.507 0.643
800 0.467 0.458 0.442 0.568 0.454 0.552 0.452 0.589
1600 0.419 0.370 0.499 0.763 0.452 0.412 0.435 0.439
3200 0.412 0.289 0.383 0.540 0.289 0.322 0.395 0.335
3 400 1.124 0.677 1.068 0.882 0.686 0.554 0.935 0.828
800 0.713 0.563 0.685 0.630 0.585 0.446 1.059 0.746
1600 0.712 0.465 0.861 0.529 0.498 0.369 0.746 0.565
3200 0.765 0.391 0.468 0.397 0.385 0.261 0.528 0.416
Appendix K mDR-learner - Panel data
K.1 Setting
@@
where $G(0)=1$ and $G(j)=\prod_{i=1}^{j}P(C_{i}=1|A=1,\bar{\textbf{C}}_{i-1}=\textbf{1},\bar{\textbf{Z}}_{i-1})$ for $j\in(1,...,N)$.

### v2

@@ Appendix D Derivation of the EIF used within the mDR-learner and mEP-learner
$\|\hat{\theta}-\theta^{*}\|_{2}^{2}\leq$ $\frac{1}{\frac{\lambda}{2}-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]+$
$\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|_{2}^{4}+\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{\mu}^{0}-\mu^{0}_{0}\|_{2}^{2}+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{G}-G_{0}\|_{2}^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}^{2}\|\hat{G}-G_{0}\|_{2}^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|_{2}^{4})$
@@
$\|\hat{\theta}-\theta^{*}\|_{2}^{2}\leq$ $\frac{1}{\frac{\lambda}{2}-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})+\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|_{2}^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{\mu}^{0}-\mu^{0}_{0}\|_{2}^{2}+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{G}-G_{0}\|_{2}^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}^{2}\|\hat{G}-G_{0}\|_{2}^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|_{2}^{4})$
@@
$L(\hat{\theta},g_{0})-L(\theta^{*},g_{0})\leq$ $D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]+\frac{\beta_{1}}{2}\|\hat{\theta}-\theta^{*}\|_{2}^{2}$
$\leq$ $\frac{\beta_{1}}{\frac{\lambda}{2}-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})+\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|_{2}^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{\mu}^{0}-\mu^{0}_{0}\|_{2}^{2}+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{G}-G_{0}\|_{2}^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}^{2}\|\hat{G}-G_{0}\|_{2}^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|_{2}^{4})-$
@@
$L(\hat{\theta},g_{0})-L(\theta^{*},g_{0})\leq$ $\frac{\beta_{1}}{\frac{\lambda}{2}-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})+\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|_{2}^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{\mu}^{0}-\mu^{0}_{0}\|_{2}^{2}+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{G}-G_{0}\|_{2}^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}^{2}\|\hat{G}-G_{0}\|_{2}^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|_{2}^{4})$
@@
Appendix G Half-sample bootstrap confidence intervals
Obtaining confidence intervals (CIs) for CATE estimates which are generated using non-parametric estimation techniques is challenging, with the required theoretical convergence guarantees typically not met in these settings. However, recent work by Ritzwoller and Syrgkanis [2024] presents a solution which can estimate CIs when the CATE is both estimated using a kernel based pseudo-outcome regression and the pseudo-outcomes used in that regression take the following form:
$\mu^{1}(Z)-\mu^{0}(Z)+\beta(A,Z)(Y-\mu^{A}(Z)).$
In the context of our work, the pseudo-outcomes from the DR-learner take this from when $\beta(A,Z)=\frac{\left(A-\pi(Z)\right)}{\pi(Z)\left(1-\pi(Z)\right)}$ and the pseudo-outcomes from the mDR-learner take this form when $\beta(A,Z)=\frac{\left(A-\pi(Z)\right)C}{\pi(Z)\left(1-\pi(Z)\right)G(A,Z)}$. The half-sample bootstrap approach then estimates CIs by approximating the distribution of
$R(X)=\hat{\theta}(X)-\theta(X),$
where $\hat{\theta}(X)$ is the CATE estimate obtained using the full sample, and $\theta(X)$ is the true CATE. It does so using the conditional distribution of
$R^{*}(X)=\hat{\theta}_{h}(X)-\hat{\theta}(X)$
where $\hat{\theta}_{h}(X)$ is the CATE estimate obtained using a half sample of the data. Ritzwoller and Syrgkanis [2024] then shows that these $R^{*}$’s can then be used to estimate $\hat{\lambda}^{2}(X)$, the variance of $\sqrt{n}R^{*}(X)$, and $\hat{cv}(\alpha)$, which denotes the $1-\alpha$ quantile of the distribution of the studentised statistic $\hat{S}^{*}(X)=\sqrt{n}\left\|\hat{\Lambda}^{-1/2}R^{*}(X)\right\|_{\infty}$, where $\hat{\Lambda}$ denotes the diagonal matrix with elements $\hat{\lambda}^{2}$. By doing so, CIs can then be constructed as:
$\hat{C}(X)=\hat{\theta}(X)\pm n^{-1/2}\hat{\lambda}(X)\hat{cv}(\alpha)$
This approach does not require the nuisance models to be re-run, however, for each bootstrap half-sample, the DR-learner’s/mDR-learner’s second stage pseudo-outcome regression must be re-run to obtain half sample estimates. We do this using a random forest, inline with the the approach used in the original work. Additionally, we implement this approach for our variations of the EP-learner, noting that the targeting step must also be re-run for each half sample bootstrap. We evaluate the performance of this approach for each of our estimators by implementing it within our simulation study. We run 500 bootstrap samples to generate the CIs and review the distribution of conditional coverages obtained within the test data. Results from this analysis can be found in Appendix L, along with a short summary of these findings.
In addition, we implement this approach on the GBSG2 data example. However, we note two key limitations of using this technique in practice. Firstly, the implementation of this approach was only tested by Ritzwoller and Syrgkanis [2024] using untuned random forests. This means the point estimates obtained when using this approach are likely to be worse than the point estimates obtained by the SuperLearner approach presented in Section 5 in the main text. Equally, CIs cannot be provided for the estimates obtained using the median estimator approach, as this theory does not extend to estimators constructed in this way. For these reasons, we report the findings of this additional GBSG2 analysis in Appendix M and only provide CIs for estimates obtained from one iteration of each learner.
Appendix H Simulation study DGP specifications
@@
$Y(0)|Z\sim\mu_{0}(Z)+\epsilon,~{}~{}~{}Y(1)|Z,Y(0)\sim Y(0)+\theta(Z),~{}~{}~{}\epsilon|Z\sim\mathcal{N}(0,1)$
@@
In each DGP, 6 covariates were created, each drawn from a uniform distribution ranging from -1 to 1. The propensity score, $\pi(Z)$, used within each DGP is defined as $\pi(Z)=expit(\sum_{i=1}^{3}(Z_{i}^{2}-sin(3Z_{i}))/1.5)$, and each DGP uses a combination of the following censoring, unexposed outcome and CATE functions. The complex unexposed outcome function was defined as $\mu^{0}(Z)=\sum_{i=3}^{4}(Z_{i}/2+4\mathbbm{1}\{Z_{i}>0.5\}+sin(4Z_{i}))/2$, while the simple unexposed outcome function was defined as $\mu^{0}(Z)=Z_{3}+Z_{4}$. The complex CATE was defined as $\theta(Z)=\mathbbm{1}\{|Z_{3}|>0.5\}+\mathbbm{1}\{|Z_{4}|>0.5\}$, while the simple CATE was defined as $\theta(Z)=0$. Then two censoring functions $G(A,Z)=expit(3-2.75A\sum_{i=4}^{5}\mathbbm{1}(|Z_{i}|>0.5))$ and $G(Z)=expit(3-2.75\sum_{\{4,5\}}\mathbbm{1}(|Z_{i}|>0.5))$ were defined, with the first imposing strong censoring only on the exposed arm, and the second imposing strong censoring evenly across both arms.
@@ 2 Complex Complex (Does not depend on A) Simple Complex
3 Complex Complex (Depends on A) Complex Simple
Appendix I Super Learner algorithm inclusion
Table 3: Simulation study - Super learner algorithm libraries
@@ Appendix I Super Learner algorithm inclusion
Table 4: GBSG2 analysis - Super learner algorithm libraries
Algorithm Tuning parameters
Mean (SL.mean) .
Linear model/Generalised linear model (SL.lm/SL.glm) .
LASSO/Elastic net (SL.glmnet) nlambda = (100,250)
alpha = (0.5,1)
Random forest (SL.ranger) mtry = (1,2)
min.node.size = (10,20,50)
Neural nets (SL.nnet) size = (1,2,5)
SVM (SL.svm) nu = 1
type = C-classification
Kernel KNN (SL.kernelKnn) h = (0.05,0.25)
Neural nets, SVMs and Kernel KNN was only used in pseudo-outcome regressions
Appendix J Root mean square median error (RMSME)
As the DR-learner and mDR-learner use IPTWs/IPCWs, extreme propensity score/censoring probability estimates can lead to unstable CATE estimates for certain individuals. Within the simulation study, this only occurs within a small subset of the simulations and consequently, the distribution of RMSE estimates ends up skewed. For this reason, a metric such as the mean RMSE can poorly represent the performance of these estimators.
Alternatively, we compare the estimators within the simulation study using the root mean square median error (RMSME). We calculate this metric by identifying the error in individual $j$’s CATE estimate for simulation $i$, $\epsilon_{ij}$. We then take the median of these errors across the simulations, $\epsilon_{j}^{med}=Median(\epsilon_{i,j})$. By doing so, extreme CATE estimates for each individual will become less influential and we can then estimate the root mean squared error (RMSE) using the median errors. This metric has a similar principle to the median CATE implementation option that was outlined in Section 3.3, reducing the estimators sensitivity to extreme CATE estimates and providing a metric that ignores outlier estimates.
@@
Appendix K Additional simulation study results
@@
DGP N mDR mEP DR EP T IPTW-IPCW
AC Imputation AC Imputation AC Imputation
1 400 0.480 0.417 0.497 0.537 0.436 0.507 0.385 0.477 0.529
800 0.360 0.314 0.376 0.410 0.329 0.379 0.318 0.376 0.486
1600 0.258 0.236 0.273 0.305 0.249 0.289 0.258 0.307 0.493
3200 0.169 0.164 0.188 0.219 0.178 0.207 0.207 0.250 0.557
2 400 0.471 0.425 0.473 0.555 0.429 0.593 0.379 0.545 0.780
800 0.354 0.325 0.347 0.422 0.325 0.449 0.317 0.428 0.890
1600 0.247 0.242 0.246 0.309 0.244 0.335 0.255 0.340 1.129
3200 0.162 0.182 0.169 0.222 0.185 0.244 0.206 0.280 1.562
3 400 0.317 0.329 0.406 0.313 0.416 0.274 0.486 0.303 2.457
800 0.236 0.265 0.331 0.261 0.360 0.236 0.417 0.264 2.360
1600 0.197 0.236 0.289 0.241 0.325 0.231 0.346 0.254 2.506
3200 0.158 0.206 0.261 0.219 0.293 0.231 0.300 0.215 2.902
DGP = Data generating process; N = Training sample size; AC = Available case; RMSME = Root mean squared median error.
Figure 8: Mean RMSE for mDR-learner, mEP-learner, DR-learner, EP-learner and T-learner in three DGPs plotted by training sample size. Plots in the left column compare the mDR-learner and mEP-learner to the DR-learner, EP-learner and T-learner when used in combination with an outcome imputation model in DGP 1, 2 and 3 respectively. Plots in the right column compare the mDR-learner and mEP-learner to the available case versions of the DR-learner, EP-learner and T-learner in DGP 1, 2 and 3 respectively.
@@
DGP N mDR mEP DR EP T IPTW-IPCW
AC Imputation AC Imputation AC Imputation
1 400 0.520 0.484 0.469 0.529 0.565 0.552 0.571 0.608 0.659
800 0.407 0.374 0.362 0.421 0.444 0.424 0.464 0.488 0.606
1600 0.323 0.298 0.294 0.320 0.341 0.332 0.415 0.363 0.609
3200 0.253 0.240 0.235 0.238 0.253 0.246 0.466 0.304 0.656
2 400 0.528 0.488 0.486 0.516 0.588 0.631 0.507 0.671 1.059
800 0.421 0.394 0.379 0.395 0.462 0.504 0.452 0.601 1.150
1600 0.329 0.315 0.297 0.294 0.346 0.376 0.435 0.442 1.388
3200 0.290 0.232 0.258 0.235 0.263 0.295 0.395 0.415 1.867
3 400 0.547 0.646 0.597 0.574 0.472 0.516 0.935 0.881 2.865
800 0.423 0.543 0.468 0.464 0.380 0.425 1.059 0.775 2.727
1600 0.315 0.452 0.372 0.417 0.341 0.386 0.746 0.758 2.845
3200 0.255 0.374 0.318 0.335 0.275 0.305 0.528 0.420 3.245
DGP = Data generating process; N = Training sample size; AC = Available case; RMSE = Root mean squared error.
Appendix L 95% confidence interval coverage results
The performance of the half sample bootstrap procedure varied across our simulation settings. In settings with a complex CATE (DGP 1 and 2), the marginal coverage increased as sample size grew, while for settings with a simple CATE (DGP 3), the procedure provided good coverage in lower sample sizes, but this coverage dropped as the sample size increased. We also note that for settings with a complex CATE, the conditional coverage for some individuals was very poor, with the minimum 95% CI coverage in single digits. Meanwhile, in settings where the CATE was simple, the minimum conditional coverage was higher, but did begin to fall as the sample size increased. Finally, we note that all variations of our two estimators performed consistently and similar trends in CI coverage were seen for all estimators.
Table 7: Summary of conditional 95% confidence interval coverage for
each learner by sample size - Example 1
Learner Sample size Mean Median SD Range
% % (Min, Max)
mDR 400 79.0 95.8 24.1 (3.0, 100.0)
800 91.2 98.8 15.1 (4.4, 100.0)
1600 95.7 100.0 10.2 (10.6, 100.0)
3200 98.0 100.0 7.1 (16.8, 100.0)
mEP 400 87.0 98.4 20.4 (2.0, 100.0)
800 90.2 99.8 18.4 (1.8, 100.0)
1600 93.2 99.8 14.5 (3.2, 100.0)
3200 96.1 99.8 9.2 (6.8, 100.0)
DR - AC 400 69.7 87.2 31.8 (1.2, 100.0)
800 84.1 98.0 24.6 (1.2, 100.0)
1600 90.4 99.8 19.2 (3.2, 100.0)
3200 95.7 100.0 11.7 (5.2, 100.0)
DR - Imputation 400 64.2 75.7 34.2 (0.4, 100.0)
800 80.1 95.8 28.5 (1.0, 100.0)
1600 87.2 99.6 23.7 (0.8, 100.0)
3200 92.7 100.0 16.8 (3.0, 100.0)
EP - AC 400 84.4 97.2 23.6 (2.6, 100.0)
800 88.5 99.8 21.4 (2.4, 100.0)
1600 91.5 99.6 17.4 (2.2, 100.0)
3200 94.2 99.8 12.2 (4.4, 100.0)
EP - Imputation 400 80.4 94.0 26.5 (2.0, 100.0)
800 87.4 99.4 22.5 (4.0, 100.0)
1600 91.4 99.6 17.7 (2.6, 100.0)
3200 95.0 99.8 12.3 (5.2, 100.0)
AC = Available case; SD = Standard deviation across conditional coverage percentages
Table 8: Summary of conditional 95% confidence interval coverage for
each learner by sample size - Example 2
Learner Sample size Mean Median SD Range
% % (Min, Max)
mDR 400 83.5 97.0 19.7 (7.4, 100.0)
800 92.5 98.8 12.6 (11.0, 100.0)
1600 97.2 100.0 7.1 (16.2, 100.0)
3200 98.9 100.0 4.6 (25.2, 100.0)
mEP 400 83.0 96.4 23.6 (3.4, 100.0)
800 86.6 99.4 22.5 (3.4, 100.0)
1600 88.8 99.0 18.8 (3.2, 100.0)
3200 90.0 99.0 16.9 (3.8, 100.0)
DR - AC 400 70.1 83.2 30.8 (1.4, 100.0)
800 82.8 98.2 26.0 (2.2, 100.0)
1600 88.5 99.4 19.9 (3.8, 100.0)
3200 93.1 99.4 13.0 (4.2, 100.0)
DR - Imputation 400 60.9 71.4 32.6 (0.8, 100.0)
800 74.6 94.0 32.8 (0.2, 100.0)
1600 80.1 97.6 29.1 (0.8, 100.0)
3200 84.9 99.0 23.1 (2.0. 100.0)
EP - AC 400 81.2 96.2 25.0 (2.8, 100.0)
800 85.8 99.2 23.6 (2.6, 100.0)
1600 87.8 98.8 20.4 (3.4, 100.0)
3200 88.1 98.6 19.8 (3.2, 100.0)
EP - Imputation 400 69.3 85.8 32.2 (1.2, 100.0)
800 77.4 97.0 30.9 (2.0, 100.0)
1600 81.2 97.0 26.8 (0.6, 100.0)
3200 86.2 98.6 21.2 (3.0, 100.0)
AC = Available case; SD = Standard deviation across conditional coverage percentages
Table 9: Summary of conditional 95% confidence interval coverage for
each learner by sample size - Example 3
Learner Sample size Mean Median SD Range
% % (Min, Max)
mDR 400 99.5 99.8 0.8 (93.4, 100.0)
800 99.5 99.8 0.9 (91.4, 100.0)
1600 99.2 99.8 1.8 (81.2, 100.0)
3200 98.9 100.0 2.9 (72.0, 100.0)
mEP 400 98.3 98.8 1.6 (90.6, 100.0)
800 98.5 99.4 2.2 (85.6, 100.0)
1600 97.2 99.4 4.6 (72.6, 100.0)
3200 95.9 99.8 8.0 (55.8, 100.0)
DR - AC 400 96.9 98.2 3.8 (77.6, 100.0)
800 96.6 98.8 5.8 (64.6, 100.0)
1600 94.8 99.2 9.6 (38.0, 100.0)
3200 93.7 99.8 13.7 (17.6, 100.0)
DR - Imputation 400 98.6 99.2 1.9 (88.0, 100.0)
800 98.4 99.6 3.5 (73.4, 100.0)
1600 96.9 99.8 7.1 (45.4, 100.0)
3200 95.3 99.8 11.3 (16.8, 100.0)
EP - AC 400 97.5 98.4 2.9 (84.0, 100.0)
800 96.9 99.2 5.2 (70.6, 100.0)
1600 94.3 99.0 9.3 (48.2, 100.0)
3200 92.1 99.6 14.7 (20.0, 100.0)
EP - Imputation 400 99.8 100.0 0.3 (97.8, 100.0)
800 99.8 100.0 0.6 (95.4, 100.0)
1600 99.1 100.0 2.1 (84.6, 100.0)
3200 98.2 100.0 4.6 (58.4, 100.0)
AC = Available case; SD = Standard deviation across conditional coverage percentages
Appendix M GBSG2 analysis using random forests for pseudo-outcome regressions
Figure 9: Median CATE estimates plotted by progesterone receptor (fmol/l).
Figure 10: CATE estimates from single cross-fitting seeds plotted by progesterone receptor (fmol/l).
Appendix N mDR-learner - Longitudinal data
N.1 Setting
@@
where $G(0)=1$ and $G(j)=\prod_{i=1}^{j}P(C_{i}=1|A=1,\bar{\textbf{C}}_{i-1}=\textbf{1},\bar{\textbf{Z}}_{i-1})$ for $j\in(1,...,L)$.